A man is 3 times as old as his daughter. 4 years ago, the sum of their ages was 60 years. How old is the daughter?

Answer by stanbon(75887)

A man is 3 times as old as his daughter.
4 years ago, the sum of their ages was 60 years.
How old is the daughter?
-----------------
Equations:
f = 3d
(f-4)+(d-4) = 60
--------------------
Substitute for "f" and solve for "d":
(3d-4)+(d-4) = 60
4d-8 = 60
4d = 68
d = 17 (daughter's age now)
---
Solve for "f":
f = 3d
f = 3*17 = 51 (father's age now)
=================================
Cheers,
Stan H.
----